# The adjacent Hindman’s theorem and the $\mathbb{Z}$-Ramsey’s theorem

### v2

@@ 1. Introduction
Many of the principles mentioned in the previous paragraph turn out to be equivalent to $\mathnormal{\mathsf{ACA}}_{0}$ over $\mathnormal{\mathsf{RCA}}_{0}$, so for the most part the viewpoint of Reverse Mathematics will not be very informative in that respect. In view of this, we focus more on the Computability-theoretic aspect of the principles, thinking instead about the reducibility relation, by means of which an instance of one problem can be turned into an instance of a different problem and a solution to the latter gets turned back into a solution to the original instance. More concretely, we will work with Weihrauch reducibility [^1] Some authors (e.g., [11, p. 22]) have called this notion uniform reducibility, but the name “Weihrauch reducibility” seems to be the most commonly used nowadays. We are grateful to two anonymous referees for this observation.. Recall [8, Definition 4.3.1] that a problem $P$ is Weihrauch reducible to another problem $Q$, denoted $P\mathop{\leq_{\mathsf{W}}}Q$, if there are Turing functionals $\Phi,\Psi$ such that for every instance $X$ of $P$, $\Phi^{X}$ is an instance of $Q$, and for every solution $Y$ to $\Phi^{X}$, $\Psi^{X\oplus Y}$ is a solution to $X$. If we have the exact same situation, but with a Turing functional $\Psi$ such that $\Psi^{Y}$ is a solution to $X$ (i.e. we may compute the solution to $X$ just with information from $Y$ without needing to use any information about $X$), then we say that $P$ is strongly Weihrauch reducible [8, Definition 4.4.1] to $Q$ and write $P\mathop{\leq_{\mathrm{s}\mathsf{W}}}Q$. If each of the problems $P,Q$ is Weihrauch reducible (respectively, strongly Weihrauch reducible) to the other, then we say that $P$ is Weihrauch equivalent (respectively, strongly Weihrauch equivalent) to $Q$ and we denote this by $P\mathop{\equiv_{\mathsf{W}}}Q$ (respectively, $P\mathop{\equiv_{\mathrm{s}\mathsf{W}}}Q$). So we will use this notion of reducibility to gauge the position of the Adjacent Hindman’s Theorem and the $\mathbb{Z}$-Ramsey’s Theorem within the computability-theoretic hierarchy; an important feature of all of our proofs is that they can be formalized within $\mathnormal{\mathsf{RCA}}_{0}$, thus yielding corresponding (but coarser) Reverse Mathematics results as corollaries.
@@ 2. The $\mathbb{Z}$-invariant Ramsey’s theorem
It is clear that, for each $n\geq 1$, -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n}$ is a trivial corollary of $\mathnormal{\mathsf{RT}}^{n}$. Now, in the case of $\mathnormal{\mathsf{RT}}^{n}$, there is no difference (whether from the viewpoint of Weihrauch equivalence, or over $\mathnormal{\mathsf{RCA}}_{0}$) whether one formulates it over $\mathbb{N}$ or over any infinite $X\subseteq\mathbb{N}$ [8, Proposition 8.3.4]. This is not the case for -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n}$: for example, if $X\subseteq\mathbb{N}$ is any infinite set with the property that (if we enumerate its elements increasingly by $x_{n}$) $x_{n+1}-x_{n}>x_{n}-x_{n-1}$ for all $n$, then (since no two distinct pairs of elements of $X$ are at the same distance from one another) the family of $\mathbb{Z}$-invariant colourings restricted to $X$ is the same as the family of all arbitrary colourings on $X$. Therefore, for such a set, the version of -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n}$ stated for $X$ would be Weihrauch-equivalent to the full $\mathnormal{\mathsf{RT}}^{n}$. We therefore refrain from considering “general” versions of the $\mathbb{Z}$-Ramsey theorem.
@@ 3.1. The apartness condition
We summarize the results of the paper, from the perspective of Reverse Mathematics, by presenting the diagram in Figure 1. Whether any of the remaining arrows (except for the one connecting $\mathnormal{\mathsf{ACA}}_{0}$ to $\mathnormal{\mathsf{RT}}^{2}$) is reversible remains an open question. For example, it is conceivable that $\mathnormal{\mathsf{AHT}}$ already implies $\mathnormal{\mathrm{ap}\mathsf{AHT}}$, although some ideas that have been used for similar results (e.g. [10, Lemma 1]) do not seem to work in this context because the requirement that the sums be adjacent, thus preventing us to skip summands, poses a strong restriction. Similar difficulties arise when one attempts to obtain better lower bounds for -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{2}$: Carlucci [2, Prop. 3] has found that the Increasing Polarized Ramsey’s Theorem for pairs [^4] That is, the statement that for every finite colouring $c:[\mathbb{N}]^{2}\longrightarrow k$ there are two infinite sets $X,Y$ such that all pairs of the form $\{x,y\}$ with $x\in X$, $y\in Y$, and $x<y$, have the same colour., denoted $\mathsf{IPT}^{2}$, is a lower bound for $\mathnormal{\mathrm{ap}\mathsf{AHT}}$; the apartness condition plays a crucial rôle in Carlucci’s proof, and we have been unable to successfully find a suitable lower bound for $\mathnormal{\mathsf{AHT}}$ without this condition [^5] It is conceivable that some of the ideas from [5] could be useful to obtain such lower bound results without separation/apartness. We are grateful to one of the anonymous referees for this suggestion.. We do have a conjecture but state it in the form of a question below. A final consideration arises from thinking about versions of the principles we have stated for a bounded number of colours. For example, one could consider, e.g., the principle $\mathnormal{\mathsf{AHT}}_{k}$ which states the same as $\mathnormal{\mathsf{AHT}}$ but only for colourings with at most $k$ colours. Note that, from this viewpoint, Theorem 14 really proves that -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{3}_{4}$ implies $\mathnormal{\mathsf{ACA}}_{0}$, but it is not clear whether that $4$ could be lowered even further. We finalize the paper by stating explicitly these questions.
@@
The principle $\mathsf{D}^{2}$ mentioned in point (2) of Questions 15 is the statement, as considered in [6], that for every stable $c:[\mathbb{N}]^{2}\longrightarrow k$ (where “stable” means for each $n$, the function $m\longmapsto c(\{n,m\})$ is eventually constant) there exists an infinite set $X\subseteq\mathbb{N}$ and a colour $i<k$ such that for each $n\in X$, we have ${\lim_{m\to\infty}c(\{n,m\})=i}$; this statement is implied by $\mathsf{IPT}^{2}$ by a result of Dzhafarov and Hirst [6]. On the other hand, regarding point (3) of Questions 15, it is worth noting that D. Tavernelli [17, Section 2.3] has obtained results showing that the versions of $\mathnormal{\mathsf{AHT}}$ with more colours are not Weihrauch-reducible to the versions with fewer colours, although this still leaves open the possibility of having some version of $\mathnormal{\mathsf{AHT}}$ with more colours being provable (over $\mathnormal{\mathsf{RCA}}_{0}$, or maybe over $\mathnormal{\mathsf{RCA}}_{0}+\mathsf{B}\Sigma_{2}^{0}$) from another version with fewer colours.

### v3

@@ 1. Introduction
Many of the principles mentioned in the previous paragraph turn out to be equivalent to $\mathnormal{\mathsf{ACA}}_{0}$ over $\mathnormal{\mathsf{RCA}}_{0}$, so for the most part the viewpoint of Reverse Mathematics will not be very informative in that respect. In view of this, we focus more on the Computability-theoretic aspect of the principles, thinking instead about the reducibility relation, by means of which an instance of one problem can be turned into an instance of a different problem and a solution to the latter gets turned back into a solution to the original instance. More concretely, we will work with Weihrauch reducibility [^1] Some authors (e.g., [9, p. 22]) have called this notion uniform reducibility, but the name “Weihrauch reducibility” seems to be the most commonly used nowadays. We are grateful to two anonymous referees for this observation.. Recall [7, Definition 4.3.1] that a problem $P$ is Weihrauch reducible to another problem $Q$, denoted $P\mathop{\leq_{\mathsf{W}}}Q$, if there are Turing functionals $\Phi,\Psi$ such that for every instance $X$ of $P$, $\Phi^{X}$ is an instance of $Q$, and for every solution $Y$ to $\Phi^{X}$, $\Psi^{X\oplus Y}$ is a solution to $X$. If we have the exact same situation, but with a Turing functional $\Psi$ such that $\Psi^{Y}$ is a solution to $X$ (i.e. we may compute the solution to $X$ just with information from $Y$ without needing to use any information about $X$), then we say that $P$ is strongly Weihrauch reducible [7, Definition 4.4.1] to $Q$ and write $P\mathop{\leq_{\mathrm{s}\mathsf{W}}}Q$. If each of the problems $P,Q$ is Weihrauch reducible (respectively, strongly Weihrauch reducible) to the other, then we say that $P$ is Weihrauch equivalent (respectively, strongly Weihrauch equivalent) to $Q$ and we denote this by $P\mathop{\equiv_{\mathsf{W}}}Q$ (respectively, $P\mathop{\equiv_{\mathrm{s}\mathsf{W}}}Q$). So we will use this notion of reducibility to gauge the position of the Adjacent Hindman’s Theorem and the $\mathbb{Z}$-Ramsey’s Theorem within the computability-theoretic hierarchy; an important feature of all of our proofs is that they can be formalized within $\mathnormal{\mathsf{RCA}}_{0}$, thus yielding corresponding (but coarser) Reverse Mathematics results as corollaries.
@@ 2. The $\mathbb{Z}$-invariant Ramsey’s theorem
It is clear that, for each $n\geq 1$, -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n}$ is a trivial corollary of $\mathnormal{\mathsf{RT}}^{n}$. Now, in the case of $\mathnormal{\mathsf{RT}}^{n}$, there is no difference (whether from the viewpoint of Weihrauch equivalence, or over $\mathnormal{\mathsf{RCA}}_{0}$) whether one formulates it over $\mathbb{N}$ or over any infinite $X\subseteq\mathbb{N}$ [7, Proposition 8.3.4]. This is not the case for -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n}$: for example, if $X\subseteq\mathbb{N}$ is any infinite set with the property that (if we enumerate its elements increasingly by $x_{n}$) $x_{n+1}-x_{n}>x_{n}-x_{0}$ for all $n$, then (since no two distinct pairs of elements of $X$ are at the same distance from one another) the family of $\mathbb{Z}$-invariant colourings restricted to $X$ is the same as the family of all arbitrary colourings on $X$. Therefore, for such a set, the version of -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n}$ stated for $X$ would be Weihrauch-equivalent to the full $\mathnormal{\mathsf{RT}}^{n}$. We therefore refrain from considering “general” versions of the $\mathbb{Z}$-Ramsey theorem.
@@ 3.1. The apartness condition
We summarize the results of the paper, from the perspective of Reverse Mathematics, by presenting the diagram in Figure 1. Whether any of the remaining arrows (except for the one connecting $\mathnormal{\mathsf{ACA}}_{0}$ to $\mathnormal{\mathsf{RT}}^{2}$) is reversible remains an open question. For example, it is conceivable that $\mathnormal{\mathsf{AHT}}$ already implies $\mathnormal{\mathrm{ap}\mathsf{AHT}}$, although some ideas that have been used for similar results (e.g. [8, Lemma 1]) do not seem to work in this context because the requirement that the sums be adjacent, thus preventing us to skip summands, poses a strong restriction. Similar difficulties arise when one attempts to obtain better lower bounds for -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{2}$: Carlucci [2, Prop. 3] has found that the Increasing Polarized Ramsey’s Theorem for pairs [^4] That is, the statement that for every finite colouring $c:[\mathbb{N}]^{2}\longrightarrow k$ there are two infinite sets $X,Y$ such that all pairs of the form $\{x,y\}$ with $x\in X$, $y\in Y$, and $x<y$, have the same colour., denoted $\mathsf{IPT}^{2}$, is a lower bound for $\mathnormal{\mathrm{ap}\mathsf{AHT}}$; the apartness condition plays a crucial rôle in Carlucci’s proof, and we have been unable to successfully find a suitable lower bound for $\mathnormal{\mathsf{AHT}}$ without this condition [^5] It is conceivable that some of the ideas from [5] could be useful to obtain such lower bound results without separation/apartness. We are grateful to one of the anonymous referees for this suggestion.. We do have a conjecture but state it in the form of a question below. A final consideration arises from thinking about versions of the principles we have stated for a bounded number of colours. For example, one could consider, e.g., the principle $\mathnormal{\mathsf{AHT}}_{k}$ which states the same as $\mathnormal{\mathsf{AHT}}$ but only for colourings with at most $k$ colours. Note that, from this viewpoint, Theorem 14 really proves that -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{3}_{4}$ implies $\mathnormal{\mathsf{ACA}}_{0}$, but it is not clear whether that $4$ could be lowered even further. We finalize the paper by stating explicitly these questions.
@@
The principle $\mathsf{D}^{2}$ mentioned in point (2) of Questions 15 is the statement, as considered in [6], that for every stable $c:[\mathbb{N}]^{2}\longrightarrow k$ (where “stable” means for each $n$, the function $m\longmapsto c(\{n,m\})$ is eventually constant) there exists an infinite set $X\subseteq\mathbb{N}$ and a colour $i<k$ such that for each $n\in X$, we have ${\lim_{m\to\infty}c(\{n,m\})=i}$; this statement is implied by $\mathsf{IPT}^{2}$ by a result of Dzhafarov and Hirst [6]. On the other hand, regarding point (3) of Questions 15, it is worth noting that D. Tavernelli [15, Section 2.3] has obtained results showing that the versions of $\mathnormal{\mathsf{AHT}}$ with more colours are not Weihrauch-reducible to the versions with fewer colours, although this still leaves open the possibility of having some version of $\mathnormal{\mathsf{AHT}}$ with more colours being provable (over $\mathnormal{\mathsf{RCA}}_{0}$, or maybe over $\mathnormal{\mathsf{RCA}}_{0}+\mathsf{B}\Sigma_{2}^{0}$) from another version with fewer colours.